OpenStudy (anonymous):

Which sequence is geometric and has 14 as its fifth term and 12 as the common ratio? …, 1, 1/2, 1/4, 1/8, … …, 1/4, 1/2, 1, 2,… …, 1/72, 1/36,1/18, 1/9, … …, 8, 4, 2, 1, …

OpenStudy (anonymous):

none of these

OpenStudy (anonymous):

unless you mean \(\frac{1}{4}\) as its fifth term and \(\frac{1}{2}\) as the common ratio

OpenStudy (anonymous):

yeah that what i mean can you please help @satellite73

OpenStudy (anonymous):

can you post a screen shot of the actual question? all your answers seem to have \(...\) at the beginning and at the end

OpenStudy (anonymous):

the first …, 1, 1/2, 1/4, 1/8, … and the last …, 1/4, 1/2, 1, 2,… both have \(\frac{1}{2}\) as the common ratio

OpenStudy (anonymous):

oops i copied that wrong the last is …, 8, 4, 2, 1, …

OpenStudy (anonymous):

Which sequence is geometric and has 1/4 as its fifth term and 1/2 as the common ratio? A ; …, 1, 1/2, 1/4, 1/8, … B; …, 1/4, 1/2, 1, 2,… C ; …, 1/72, 1/36,1/18, 1/9, … D ; …, 8, 4, 2, 1, … @satellite73

OpenStudy (anonymous):

A and D both have \(\frac{1}{2}\) as the common ratio,

OpenStudy (anonymous):

is the geometric series 1/4 as the 5th term

OpenStudy (anonymous):

i wish i could help A and D both have \(\frac{1}{2}\) as the common ratio but since you have \(...\) at the beginning of both, it is not clear what the first term is, so it is not clear what the fifth term is

OpenStudy (anonymous):

of the first term is \(8\) then the fifth term would be \(\frac{1}{2}\) as in D, but really i cannot understand how you are supposed to know what is the first term with the repeated dots at the beginning that is why i asked for a screen shot, maybe i could figure it out from there

OpenStudy (anonymous):

oh kay but thats i got it now !